OpenStudy (coconut):

What is the simpliified form of (x^2-36)/(x-6)

Parth (parthkohli):

= x + 6

OpenStudy (anonymous):

(x² - 36) = (x² - 6² ) = (x - 6) * (x + 6) * (x - 6 ) --> (x-6)² * (x+6)

OpenStudy (coconut):

my options are x – 6, with the restriction x ≠ –6 x – 6, with the restriction x ≠ 6 x + 6, with the restriction x ≠ 6 x + 6, with the restriction x ≠ –6

OpenStudy (anonymous):